[Hallc_running] [New Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Fri Apr 13 11:25:01 EDT 2018

Logentry Text:
<p><strong>CEBAF status</strong>:</p><ul><li>PSS drop in North Linac and resulting hardware problems prevented beam for most of the day shift on April 12. Only 0.9 hours of beam time.</li><li>Work on "Pass 2" separation for Hall B prevented beam for part of the swing shift on April 12. Only 4.0 hours of beam time.</li><li>We are now sharing the slit with Hall D (https://logbooks.jlab.org/entry/3559783).</li><li>On Monday, April 16, there will be beam studies for about eight hours. The beam studies will start at 8:00 a.m.. We are asked to remain locked down to allow for tuning for high beam currents.</li><li>RF recovery is scheduled for Tuesday, April 17, for about six hours. Controlled accesses will be possible during that time period.</li></ul><p><strong>Hall C status</strong>:</p><ul><li>Hall C is configured for the SIDIS experiment (E12-09-017). The HMS is setup for detecting electrons and the SHMS is setup for detecting charged pions and kaons. Details on the experimental setup a
 re available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan#Experimental_setup</li><li>The beam current is reduced significantly to keep the pTRIG1 rate below 500kHz.</li><li>Continued problems with PSU of HMS dipole during owl shift (https://logbooks.jlab.org/entry/3559847), required control accesses and repair by Steve Lassiter and Jack Segal (https://logbooks.jlab.org/entry/3559989).</li></ul><p><strong>Data taking</strong>:</p><ul><li>The detailed run plan is available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan</li><li>For kinematic setting 1, we have completed 7 out of 14 measurements with positive SHMS polarimetry. We have lost a lot of run time due to the CEBAF problems during the day shift and part of the swing shift (see CEBAF status) and the problems with the PSU of the HMS dipole during the owl shift (see Hall C status).</li><li>The data taking of the SIDIS experiment with various configuration changes between the runs requires thre
 e persons on shift. We have many open slots !
 in the shift schedule (https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-2017-2018), in particular for next week’s owl shift. We need more volunteers for shifts and have reached out to all members of the SIDIS/Kaon-LT rungroup.</li></ul><p><strong>Hall C Daily RC meeting</strong>: We will meet today at 1:30 p.m. in the conference room located on the second floor of the counting house. Remote participation will be possible via Blue Jeans (Meeting ID: 543903672). You can join via:</p><ul><li>browser: https://bluejeans.com/543903672</li><li>phone: (888) 240-2560 (US toll free), list of international numbers: https://bluejeans.com/numbers</li></ul><p>There will be no meetings at the weekend.</p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3560073
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180413/bf94580f/attachment.html>

More information about the Hallc_running mailing list